OpenStudy (anonymous):

I'm kinda stuck Solve the initial value problem \[y''+3y=0\] \[y(0)=1\] \[y'(0)=3\] \[r^2+3=0\] \[r=\sqrt3\] or \[r=-\sqrt3\] \[y=c_1e^{\sqrt3x}+c_2e{-\sqrt3x}\] \[y(0)=c_1+c_2=1\] \[y'=\sqrt3c_1e^{\sqrt3x}-\sqrt3c_2e{-\sqrt3x}\] \[y'(0)=\sqrt3c_1-\sqrt3c_2=3\]

OpenStudy (lgbasallote):

you made a fatal mistake already...

OpenStudy (lgbasallote):

\[r^2 + 3 = 0\] \[\implies r^2 = -3\]

OpenStudy (anonymous):

Oh my. It's an imaginary number

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Ok let's start over. Give me a minute

OpenStudy (lgbasallote):

sure. this isnt a pretty one

OpenStudy (anonymous):

It's not that ugly if you use Euler's identity :)

OpenStudy (lgbasallote):

everything in math is ugly :p lol

OpenStudy (anonymous):

\[y=c_1e^{i\sqrt3x}+c_2e^{-i\sqrt3x}\] \[y(0)=c_1+c_2=0\] for y' can I just take the derivative normally or do I have to do something special for the imaginary number?

OpenStudy (lgbasallote):

uhhh when you have \[r = a + bi\] aren't you supposed to use \[e^{ax} (c_1 \sin bx + c_2 \cos bx)\] or something that looks like that..i can't remember this much

OpenStudy (anonymous):

I'm not too familiar with imaginary numbers myself...let's see what google has to say

OpenStudy (anonymous):

I think you're right

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

Ok let's see. \[r_1, r_2: \alpha \pm i\beta\] in our case would this be: \[r_1=i\sqrt3\] \[r_2=-i\sqrt3\] \[y=e^x( c_1 cos\sqrt3x-c_2sin\sqrt3x)\]

OpenStudy (anonymous):

Nope. \(\alpha=0\), so \(e^{\alpha x}=1\).

OpenStudy (anonymous):

so \[y=( c_1 cos\sqrt3x-c_2sin\sqrt3x)\]

OpenStudy (anonymous):

Yeah, but you don't need (parentheses).

OpenStudy (anonymous):

THanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

How does this look? \[y(0)=c_1cos\sqrt3(0)-c_2sin\sqrt3(0)=1\]